OpenStudy (anonymous):

how do i rearrange these equations to make them easier for me to solve? a+s+t=90 s+t=a+20 9a+5s+2t=515

hartnn (hartnn):

you can plug in s+t = a+20 from 2nd equation directly into the 1st equation to get an equation only in 'a'

hartnn (hartnn):

can you plug that in ?

OpenStudy (anonymous):

yes i get a+t=110 correct?

hartnn (hartnn):

how could you get 't' also ? can you show your work ?

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):